Find the slope of a line passing through the points (11,-2) and the origin.
Answer by lwsshak3(11628)

Find the slope of a line passing through the points (11,-2) and the origin
***
slope=∆y/∆x=(-2-0)/(11-0)=-2/11(slope of given line)
Equation: y=-2x/11+b
solving for b
-2=-2*11/11+b
-2=-2+b
b=0
equation:y=-2x/11